A canister contains 20 pieces of candy: 5 strawberry flavored, 9 watermelon flavored, and 6 mint flavored. Two are selected at random. Find each probability.
P(1 strawberry and 1 mint)

Answer by stanbon(57290)

A canister contains 20 pieces of candy: 5 strawberry flavored, 9 watermelon flavored, and 6 mint flavored. Two are selected at random. Find each probability.
P(1 strawberry and 1 mint)
-----
# of strawberry/mint pairs: 5*6 = 30
---
# of random pairs: 20C2 = 190
---
P(strawberry/mint) = 30/190 = 3/19
======================================
Cheers,
stan H.